# Apparent teleportation of indistinguishable particles

### v2

@@ I.2 Observing apparent teleportation
Figure 3: The simplest (1+1) D model of heavy-ion collisions and charm production explains the idea of observing apparent teleportation. Processes resulting in $c$ and $\bar{c}$ quarks at hadronisation with flow velocities $\beta_{c}$ and $\beta_{\bar{c}}$ are depicted as AT - apparent teleportation and LD - local distinguishable. The corners cut by the red dashed lines are forbidden for the local distinguishable but allowed for the apparent teleportation processes.
After the hadronisation stage, the carriers of the $c$ and $\bar{c}$ quarks are charm and anticharm hadrons. For simplicity, we assume these are $D$ and $\bar{D}$ mesons. The free-streaming velocities of these mesons, $\beta_{D}$ and $\beta_{\bar{D}}$, are superpositions of the quark flow velocities ( $\beta_{c}$, $\beta_{\bar{c}}$ ) and a stochastic, non-flow component. The non-flow component may arise from initial momenta of quarks, their diffusion in quark-gluon plasma, hadronisation, and following decay of resonances. In Methods, we illustrate the effect of the smearing by the stochastic uncorrelated processes. We show the calculation results within the 1+1D model with the Hubble-like initial and hadronisation hypersurfaces and uncorrelated smearing included. The predictions for the 1+3D model with the blast-wave hadronisation hypersurface are also presented in Methods.
Assuming that the stochastic components of $D$ and $\bar{D}$ are uncorrelated, the desired joint probability density function, $\rho(\beta_{c},\beta_{\bar{c}})$, can be extracted from the measured $\rho(\beta_{D},\beta_{\bar{D}})$ in an almost model-independent manner. The procedure is detailed in Ref. [26].
@@
The apparent teleportation may render all transitions possible. One may speculate that at very high energy densities, the microstate-symmetric transition matrix [13] may govern the system’s evolution. This implies that $c$ and $\bar{c}$ quarks, and consequently $D$ and $\bar{D}$ hadrons, are produced independently in space-time and momentum. The joint probability density function $\rho(\beta_{D},\beta_{\bar{D}})$ factorises as:
$\rho(\beta_{D},\beta_{\bar{D}})=\rho(\beta_{D})\cdot\rho(\beta_{\bar{D}})\leavevmode\nobreak\ ,$ (2)
@@
$\rho(\mathbf{p}_{D},\mathbf{p}_{\bar{D}})=\rho(\mathbf{p}_{D})\cdot\rho(\mathbf{p}_{\bar{D}})\leavevmode\nobreak\ ,$ (3)
where $\mathbf{p}_{D}$ and $\mathbf{p}_{\bar{D}}$ are the momentum vectors of $D$ and $\bar{D}$ hadrons, respectively. Contemporary experiments can test this extreme prediction following the apparent teleportation hypothesis soon [26]. This requires measurements of sufficiently high statistics of charm and anticharm hadron pairs produced in individual collisions with a mean multiplicity of the pairs below one [26].
II Discussion
@@
The presented example of the apparent teleportation process suggests that the particle-antiparticle creation and annihilation reactions, real and virtual, are required for apparent teleportation—their rate increases with the system’s energy density. Thus, apparent teleportation is expected to be more popular in high-energy-density systems. It makes it difficult to observe experimentally. This is because the high-density systems created in laboratories typically have small spatial and temporal dimensions. Therefore, precisely measuring particle number distribution in short intervals may be challenging.
@@
We provide an example of overcoming the problem of observing the apparent teleportation. It can be done by measuring momentum correlation of charm and anticharm hadrons produced in collisions of two sufficiently heavy atomic nuclei with only one pair of charm and anti-charm quarks created. Contemporary experiments can perform the measurement; however, interpreting results on apparent teleportation will depend on modelling the time evolution of the strongly interacting system. Other experimental setups should be considered, giving together evidence of the apparent teleportation of indistinguishable particles.
@@
If it exists, the apparent teleportation may explain two puzzles about nuclear collisions at high energies: the fast equilibration of the created system of quarks and gluons and its fast hadronisation fulfilling the requirement of local colour neutrality.
Methods
1+1D Hubble-like model
Here, we present predictions for the apparent teleportation using a one-dimensional evolution with Hubble-like flow, including a simple model of hadronisation of heavy quarks. The calculations follow the approach we introduced in Ref. [26] for a spherically symmetric system.
@@ Methods
$t^{2}=z^{2}+\tau_{0}^{2}\leavevmode\nobreak\ ,$ (4)
where $\tau_{0}$ is the initial proper time at which the system is formed, and $z$ is the distance from the system’s centre. The system then undergoes a Hubble-like expansion, and the hadronisation hypersurface is taken as
$t^{2}=z^{2}+\tau_{\mathrm{HAD}}^{2}\leavevmode\nobreak\ ,$ (5)
where $\tau_{\mathrm{HAD}}$ is the hadronisation proper time. The expanding matter is assumed to be in local equilibrium corresponding to temperatures $T_{0}$ and $T_{\mathrm{HAD}}$ at proper times $\tau_{0}$ and $\tau_{\mathrm{HAD}}$, respectively. The matter density is assumed to be uniform along the hadronistation hypersurface, which is limited in $z$ as $|z|<\tau_{\mathrm{HAD}}+R_{\mathrm{MAX}}$. In the following the parameters $\tau_{\mathrm{HAD}}$ and $R_{\mathrm{MAX}}$ are fixed to 10 fm and 6 fm, being on the order of the Pb nucleus radius. Figure 4 shows a sketch illustrating the hypersurfaces and indicating its parameters.
@@
$\frac{\tau_{0}}{\tau_{\mathrm{HAD}}}=1-\frac{3}{4}\;\ln\left(\frac{T_{0}}{T_{\mathrm{HAD}}}\right)\leavevmode\nobreak\ .$ (6)
The hadronisation temperature is set to the well-known value from the lattice QCD, 150 MeV [18]. The initial temperature $T_{0}$ has to be larger than the hadronisation one. Below, we present examples calculated for $T_{0}=200$ MeV and $T_{0}=300$ MeV. The latter value is close to the initial temperature given by the Fermi-Landau initial conditions [24] at the top energy CERN SPS. It can be considered as the upper limit of $T_{0}$ at SPS. The remaining parameter of the model $\tau_{0}$ is calculated using Eq. 6. The velocity of the fluid element and (anti)charm quarks at the hadronisation is given by the Hubble flow assumption, $u=z/\tau_{\mathrm{HAD}}$.
@@
$dN/dp_{D}^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \propto\leavevmode\nobreak\ \leavevmode\nobreak\ {p^{\prime}_{D}}^{2}\leavevmode\nobreak\ \cdot\leavevmode\nobreak\ e^{-\sqrt{m^{2}_{D}+{p_{D}^{\prime}}^{2}}\leavevmode\nobreak\ /\leavevmode\nobreak\ T_{\mathrm{HAD}}}\leavevmode\nobreak\ ,$ (7)
@@
Figure 5 shows two examples ($T_{0}=200$ MeV (a) and $T_{0}=300$ MeV (b)) of the hadronisation points of charm and anti-charm quarks requiring the apparent teleportation - the past light cones of $c$ and $\bar{c}$ quarks overlap only below the initial hypersurface. We notice that the larger $T_{0}$ is, the smaller the space for separating the apparent teleportation from the local-distinguishable processes. The apparent teleportation processes can be separated up to $T_{0}\approx 400$ MeV within the model and for its parameters.
Figures 6 and 7 show the joint rapidity distributions of $D$ and $\bar{D}$ mesons for $\tau_{\mathrm{HAD}}=10$ fm for $T_{0}=200$ MeV and $T_{0}=300$ MeV, respectively. The space for separating the apparent teleportation is significantly larger for the lower value of $T_{0}$. Including statistical hadronisation significantly smears the joint rapidity distributions. In this case, separating the apparent teleportation will require correcting the 2D distributions for the smearing as discussed in Ref. [26].
@@
Figure 6: Joint rapidity distributions of $D$ and $\bar{D}$ mesons for the initial temperature $T_{0}=200$ MeV. The upper plots show the distributions for $D$ and $\bar{D}$ pairs, which can result only from the apparent teleportation. The lower plots show the results for the pairs, which can be attributed either to local-distinguishable processes or the apparent teleportation. The left plots are obtained without smearing due to statistical hadronisation, whereas the right ones include the smearing.
Figure 7: Joint rapidity distributions of $D$ and $\bar{D}$ mesons for the initial temperature $T_{0}=300$ MeV. The upper plots show the distributions for $D$ and $\bar{D}$ pairs, which can result only from the apparent teleportation. The lower plots show the results for the pairs, which can be attributed either to local-distinguishable processes or the apparent teleportation. The left plots are obtained without smearing due to statistical hadronisation, whereas the right ones include the smearing.
1+3D Blast-Wave model
A simple 1+3D model of high-energy nucleus-nucleus collisions with the properties necessary to investigate apparent teleportation is presented here. The model utilises the Blast-Wave model [30] of radial flow, which has been widely used for parametrising experimental results on hadron spectra for decades. For simplicity, only central Pb+Pb collisions are considered.
@@
    The hadronistation state, where the radial flow stops, is approximated as a sphere with radius $2R$.
@@
    $\mathbf{v(r)}=c\cdot\mathbf{r}/(2R)\leavevmode\nobreak\ ,$ (8)
@@
Neglecting the non-flow component of the final velocity, if the hadronistation points of $c$ and $\bar{c}$ are chosen independently, the opening angle $\Theta$ varies between 0 and $\pi$. The cosine of the opening angle, $\cos(\Theta)$, is uniformly distributed. This set of pairs includes both local-distinguishable pairs and those violating this condition.
At $t=0$, the collision light cone is a sphere with radius $R$ centred at $(0,0,0)$. Simultaneously, the past light cones of the hadronising $c$ and $\bar{c}$ quarks on a sphere with radius $r=aR$ ($0\leq a\leq 2$) are also spheres of radius $R$, but centred at $\mathbf{r}_{c}$ and $\mathbf{r}_{\bar{c}}$ ( $r_{c}=r_{\bar{c}}=aR$ ), respectively.
@@
$\Theta_{\text{MAX}}(aR)=180^{\circ}-\cos^{-1}(1-a^{2}/2)\leavevmode\nobreak\ .$ (9)
For $a=2$ (the maximum hadronisation surface radius), $\Theta_{\text{MAX}}=0$, while for $a\to 1$, $\Theta_{\text{MAX}}\to 120^{\circ}$. For $a=1$ and below, $\Theta_{\text{MAX}}=180^{\circ}$. A sketch illustrating this calculation is shown in Fig. 8.
The LD condition suppresses the opening angle distribution at large angles, with the effect strongest for pairs emitted from $r=2R$, where $\Theta_{\text{MAX}}=0^{\circ}$. For pairs hadronising at small radii ($r\leq R$) one gets $\Theta_{\text{MAX}}=180^{\circ}$.
To test for apparent teleportation, one can focus on pairs emitted from the outer hadronisation layer (high-momentum $D$ and $\bar{D}$ mesons). Observing pairs with $\Theta>\Theta_{\text{MAX}}$ at high momentum would indicate apparent teleportation of $c$ and $\bar{c}$ quarks.

### v4

@@ 2.2 Observing apparent teleportation
Figure 3: The simplest 1+1D model of heavy-ion collisions and charm production explains the concept of observing apparent teleportation. Processes resulting in $c$ and $\bar{c}$ quarks at hadronisation with flow velocities $\beta_{c}$ and $\beta_{\bar{c}}$ are depicted as AT - apparent teleportation and LD - local distinguishable. The corners cut by the red dashed lines are forbidden for the local distinguishable but allowed for the apparent teleportation processes.
After the hadronisation stage, the carriers of the $c$ and $\bar{c}$ quarks are charm and anticharm hadrons. For simplicity, we assume these are $D$ and $\bar{D}$ mesons. The free-streaming velocities of these mesons, $\beta_{D}$ and $\beta_{\bar{D}}$, are superpositions of the quark flow velocities ( $\beta_{c}$, $\beta_{\bar{c}}$ ) and a stochastic, non-flow component. The non-flow component may arise from initial momenta of quarks, their diffusion in quark-gluon plasma, hadronisation, and the following decay of resonances. In Methods A, we present a didactic illustration of the effect of the smearing of the flow components by the stochastic uncorrelated processes. We show the results obtained within the schematic 1+1D model with the Hubble-like initial and hadronisation hypersurfaces.
Finally, in Methods B, the 1+3D model of high-energy nucleus-nucleus collisions, with properties that allow for the prediction of experimentally measured quantities, is presented. It utilises the Blast-Wave model Schnedermann:1993ws of radial flow. The results indicate that the distribution of the opening angle between momenta of $D$ and $\bar{D}$ mesons may yield evidence for the apparent teleportation. Assuming that the stochastic components of $D$ and $\bar{D}$ are uncorrelated, this distribution can be corrected for the smearing in an almost model-independent manner. The procedure is detailed in Ref. Gazdzicki:2023niq .
@@
The apparent teleportation may render all transitions possible. One may speculate that at very high energy densities, the microstate-symmetric transition matrix Gazdzicki:2022zej may govern the system’s evolution. This implies that $c$ and $\bar{c}$ quarks, and consequently $D$ and $\bar{D}$ hadrons, are produced independently in space-time and momentum. The joint probability density function $\rho(\beta_{D},\beta_{\bar{D}})$ factorises as:
$\rho(\beta_{D},\beta_{\bar{D}})=\rho(\beta_{D})\cdot\rho(\beta_{\bar{D}})~,$ (2)
@@
$\rho(\mathbf{p}_{D},\mathbf{p}_{\bar{D}})=\rho(\mathbf{p}_{D})\cdot\rho(\mathbf{p}_{\bar{D}})~,$ (3)
where $\mathbf{p}_{D}$ and $\mathbf{p}_{\bar{D}}$ are the momentum vectors of $D$ and $\bar{D}$ hadrons, respectively. Contemporary experiments can test this extreme prediction following the apparent teleportation hypothesis soon Gazdzicki:2023niq . This requires measurements of sufficiently high statistics of charm and anticharm hadron pairs produced in individual collisions with a mean multiplicity of pairs below one Gazdzicki:2023niq .
3 Discussion
@@
The presented toy-model example of the apparent teleportation mechanism suggests that the coherent creation and annihilation of particle-antiparticle pairs is required for apparent teleportation. Thus, apparent teleportation is expected to be more popular in high-energy-density systems. It makes it difficult to observe experimentally. This is because the high-density systems created in laboratories typically have small spatial and temporal dimensions. Therefore, precisely measuring particle number distribution in short time intervals may be challenging.
@@
We provide an example of overcoming the problem of observing the apparent teleportation. It can be done by measuring momentum correlation of charm and anticharm hadrons produced in collisions of two sufficiently heavy atomic nuclei, with only one pair of charm and anticharm quarks created. Contemporary experiments can perform the measurement; however, interpreting results on apparent teleportation will depend on modelling the time evolution of the strongly interacting system.
To strengthen the conclusions on the apparent teleportation, other experimental setups should be considered. The example presented above concerns apparent teleportation driven by strong interactions. It can be straightforwardly extended from the charm-anticharm hadron correlations to the bottom-antibottom hadron correlations. The mean number of $b$ and $\bar{b}$ pairs, $\langle b\bar{b}\rangle$, produced in central Pb+Pb collisions at the top LHC energy is expected to be on the order of 10 He:2022tod . Thus, by lowering the LHC beam energy, the condition $\langle b\bar{b}\rangle\approx 1$ can be reached. Alternatively, one can consider the possibility of the measurements at the top RHIC energy by the operational sPHENIX experiment sPHENIX:HF . The corresponding measurements of bottom-antibottom hadron correlations may be a challenging task due to the low detection probability for central Pb+Pb collisions recorded in the collider mode.
The possibility of using strange-antistrange hadron correlations to search for the apparent teleportation should also be considered. The requirement of a mean number of strange and antistrange hadrons being on the order of one implies measuring their correlations in central Pb+Pb collisions at collision energies of several GeV Blume:2011sb . These are energies of the future CBM experiment at FAIR CBM:2016kpk . They are below the onset of the quark-gluon plasma creation NA49:2002pzu ; NA49:2007stj . Thus, the carriers of strangeness are different species of strange and antistrange hadrons and resonances. This brings the system closer to the system of distinguishable particles for which, by definition, the apparent teleportation does not exist. Thus, one expects that at energies below the onset of quark-gluon plasma creation, the chance to observe the apparent teleportation is reduced. This concerns the search for the apparent teleportation via correlations of strange-antistrange hadrons.
Finally, we consider the possibility of observing apparent teleportation driven by electromagnetic interactions. It seems that the simplest setup consists of an extended system of photons at energy density close to the Schwinger limit for spontaneous creation of $e^{+}-e^{-}$ pairs Schwinger:1951nm . In terms of the electric field strength, the Schwinger limit is about $1.3\cdot 10^{18}$ V/m, which corresponds to the energy density $\approx 10^{18}$ J/m3 ($\approx 10^{6}$ GeV/fm3). Existing and planned ultra-intense laser facilities, like the Extreme Light Infrastructure in Europe, are far below the Schwinger limit.
Also, ultra-peripheral heavy-ion collisions (UPCs) at the LHC may offer an opportunity to search for apparent teleportation driven by electromagnetic interactions. In such collisions, a photon interacts with a strong electromagnetic field created by a heavy ion moving with ultra-relativistic velocities. These conditions facilitate the production of $c\bar{c}$ pairs; see, for example, measurements of charmonium production in the UPCs by the LHC experiments CMS:2023snh ; ALICE:2021gpt ; LHCb:2021bfl or a recent report on $D^{0}$ production in UPCs by the CMS experiment CMS-PAS-HIN-24-003 . Thus, in principle, a measurement of $D-\bar{D}$ correlation is possible, simultaneously fulfilling the requirement of a single charm pair created in a collision. However, the probability of apparent teleportation is likely significantly lower under these conditions, and a feasibility study is necessary to determine an optimal experimental setup to observe it. Other probes (strangeness, $e^{+}-e^{-}$ pairs) could also be considered.
@@
If it exists, the apparent teleportation may explain two puzzles about strong interactions at high energy densities: the rapid equilibration of the created system of quarks and gluons and its rapid hadronisation, fulfilling the requirement of local colour neutrality.
Appendix A The 1+1D Hubble-like model
Here, we illustrate the effect of the smearing by stochastic uncorrelated processes, taking as an example the statistical hadronisation of charm and anticharm quarks. We present the results obtained within the schematic 1+1D model with the Hubble-like initial and hadronisation hypersurfaces.
@@ Appendix A The 1+1D Hubble-like model
$t^{2}=z^{2}+\tau_{0}^{2}~,$ (4)
where $\tau_{0}$ is the initial proper time the system is formed, and $z$ is the distance from the system’s centre. The system then undergoes a Hubble-like expansion, and the hadronisation hypersurface is taken as
$t^{2}=z^{2}+\tau_{\mathrm{HAD}}^{2}~,$ (5)
where $\tau_{\mathrm{HAD}}$ is the hadronisation proper time. The expanding matter is assumed to be in local equilibrium corresponding to temperatures $T_{0}$ and $T_{\mathrm{HAD}}$ at proper times $\tau_{0}$ and $\tau_{\mathrm{HAD}}$, respectively. The matter density is assumed to be uniform along the hadronisation hypersurface, which is limited in $z$ as $|z|<\tau_{\mathrm{HAD}}+R_{\mathrm{MAX}}$. In the following the parameters $\tau_{\mathrm{HAD}}$ and $R_{\mathrm{MAX}}$ are fixed to 10 fm and 6 fm, being on the order of the Pb nucleus radius. Figure 4 shows a sketch illustrating the hypersurfaces and indicating their parameters.
$\frac{\tau_{0}}{\tau_{\mathrm{HAD}}}=1-\frac{3}{4}\;\ln\left(\frac{T_{0}}{T_{\mathrm{HAD}}}\right)~.$ (6)
The hadronisation temperature is set to the well-known value from the lattice QCD, $T_{\textrm{HAD}}=$ 150 MeV Aoki:2006br . The initial temperature $T_{0}$ has to be larger than the hadronisation temperature. Below, we present examples calculated for $T_{0}=200$ MeV and $T_{0}=300$ MeV. The latter value is close to the initial temperature given by the Fermi-Landau initial conditions Gazdzicki:1998vd at the top energy CERN SPS. It can be considered as the upper limit of $T_{0}$ at SPS. The remaining parameter of the model $\tau_{0}$ is calculated using Eq. 6. The velocity of the fluid element and (anti) charm quarks at the hadronisation is given by the Hubble flow assumption, $u=z/\tau_{\mathrm{HAD}}$.
@@
$dN/dp_{D}^{\prime}~~\propto~~{p^{\prime}_{D}}^{2}~\cdot~e^{-\sqrt{m^{2}_{D}+{p_{D}^{\prime}}^{2}}~/~T_{\mathrm{HAD}}}~,$ (7)
@@
Figure 5 shows two examples ($T_{0}=200$ MeV (a) and $T_{0}=300$ MeV (b)) of the hadronisation points of charm and anti-charm quarks requiring the apparent teleportation - the past light cones of $c$ and $\bar{c}$ quarks overlap only below the initial hypersurface. We notice that the larger $T_{0}$ is, the smaller the space for separating the apparent teleportation from the local-distinguishable processes. The apparent teleportation processes can be separated up to $T_{0}\approx 400$ MeV within the model, with its other parameters set to the values given in the text.
Figures 6 and 7 show the joint rapidity ($y=tanh^{-1}(\beta)$) distributions of $D$ and $\bar{D}$ mesons for $\tau_{\mathrm{HAD}}=10$ fm for $T_{0}=200$ MeV and $T_{0}=300$ MeV, respectively. The space for separating the apparent teleportation is significantly larger for the lower value of $T_{0}$. Including statistical hadronisation significantly smears the joint rapidity distributions.
@@
Figure 6: Joint rapidity distributions of $D$ and $\bar{D}$ mesons for the initial temperature $T_{0}=200$ MeV. The plots a and b show the distributions for $D$ and $\bar{D}$ pairs, which can result only from the apparent teleportation. The plots c and d show the results for the pairs, which can be attributed either to local-distinguishable processes or the apparent teleportation. The plots a and c are obtained without smearing due to statistical hadronisation, whereas the plots b and d include the smearing.
Figure 7: Joint rapidity distributions of $D$ and $\bar{D}$ mesons for the initial temperature $T_{0}=300$ MeV. The plots a and b show the distributions for $D$ and $\bar{D}$ pairs, which can result only from the apparent teleportation. The c and d plots show the results for the pairs, which can be attributed either to local-distinguishable processes or the apparent teleportation. The plots  a and c were obtained without smearing due to statistical hadronisation, whereas the plots b and d include the smearing.
Appendix B The 1+3D Blast-Wave model
The 1+3D model of high-energy nucleus-nucleus collisions, with properties that allow for the prediction of experimentally measured quantities needed to observe the apparent teleportation, is presented here. The model utilises the Blast-Wave model Schnedermann:1993ws of radial flow, which has been widely used for parametrising experimental results on hadron spectra for decades. Here, only central Pb+Pb collisions are considered.
@@ Appendix B The 1+3D Blast-Wave model
    The hadronisation stage, where the radial flow stops, is approximated as a sphere with radius $2R$.
@@
    $\mathbf{v(r)}=c\cdot\mathbf{r}/(2R)~,$ (8)
@@
Neglecting the non-flow component of the final velocity, if the hadronisation points of $c$ and $\bar{c}$ are chosen independently, the cosine of the opening angle, $\cos(\Theta)$, is uniformly distributed. This set of pairs includes both local-distinguishable pairs and those violating this condition.
At $t=0$, the collision light cone is a sphere with radius $R$ centred at (0,0,0)(0,0,0)( 0 , 0 , 0 ). The past light cones of the hadronising $c$ and $\bar{c}$ quarks on a sphere with radius $r=aR$ ($0\leq a\leq 2$) are also spheres of radius $R$, but centred at $\mathbf{r}_{c}$ and $\mathbf{r}_{\bar{c}}$ ( $r_{c}=r_{\bar{c}}=aR$ ), respectively.
@@
$\Theta_{\text{MAX}}(aR)=\pi-\cos^{-1}(1-a^{2}/2)~.$ (9)
For $a=2$ (the maximum hadronisation surface radius), $\Theta_{\text{MAX}}=0$, while for $a\to 1$, $\Theta_{\text{MAX}}\to 2/3~\pi$. For $a=1$ and below, $\Theta_{\text{MAX}}=\pi$. A sketch illustrating this calculation is shown in Fig. 8.
The LD condition suppresses the opening angle distribution at large angles, with the effect strongest for pairs emitted from $r=2R$, where $\Theta_{\text{MAX}}=0^{\circ}$. For pairs hadronising at small radii ($r\leq R$) one gets $\Theta_{\text{MAX}}=\pi$.
One can focus on pairs emitted from the outer hadronisation layer (high-momentum) to test for apparent teleportation. Observing pairs with $\Theta>\Theta_{\text{MAX}}$ at high momentum would indicate apparent teleportation of $c$ and $\bar{c}$ quarks.
Various processes, in particular, hadronisation of (anti) charm quarks, will smear the opening angle distribution between $D$ and $\bar{D}$ mesons. Thus, separating the apparent teleportation will require correcting the opening angle distribution for the smearing as discussed in Ref. Gazdzicki:2023niq . As the smearing will be mostly due to effects that affect the emission angles of $D$ and $\bar{D}$ independently, the correction is expected to be approximately model-independent.